# Examining Differential Item Functioning (DIF) in Self-Reported Health Survey Data: Via Multilevel Modeling

### v1

@@ Abstract
Few health-related constructs or measures have received critical evaluation in terms of measurement equivalence, such as self-reported health survey data. Differential item functioning (DIF) analysis is crucial for evaluating measurement equivalence in self-reported health surveys, which are often hierarchical in structure. While traditional DIF methods rely on single-level models, multilevel models offer a more suitable alternative for analyzing such data. In this article, we highlight the advantages of multilevel modeling in DIF analysis and demonstrate how to apply the DIF framework to self-reported health survey data using multilevel models. For demonstration, we analyze DIF associated with population density on the probability to answer “Yes” to a survey question on depression and reveal that multilevel models achieve better fit and account for more variance compared to single-level models. This article is expected to increase awareness of the usefulness of multilevel modeling for DIF analysis and assist healthcare researchers and practitioners in improving the understanding of self-reported health survey data validity.
@@ 1 Introduction
To examine the disproportionate impact of health interventions or other determinants on health outcomes for certain populations, including disadvantaged minorities, it is crucial to establish measurement equivalence when evaluating outcomes for different demographic groups. However, few health constructs or measures have undergone critical evaluation in this regard [1], despite the routine practice of validating psychometric instruments in research. Recently, more health studies have employed differential item functioning (DIF) analysis to examine measurement equivalence and the differential impact of groups differing in gender and ethnicity [e.g., 2, 3, 4, 5, 6]. DIF analysis assesses whether there is a systematic difference in the probability of obtaining a targeted response between groups while controlling for a latent trait [7]. That said, some health studies did not clearly describe the method used for DIF analysis [e.g., 3], and conventional DIF methods based on single-level models were frequently used to analyze health survey data with a nesting structure [e.g., 2], for which recent DIF methods based on multilevel models are better alternatives in terms of accuracy of DIF estimation.
@@ 2 Challenge: Examine DIF in self-reported health survey data
Existing research has shown that bias can manifest itself as construct bias, method bias, and item bias [e.g., 11, 12]. The framework of differential item functioning [DIF; 13] underpins many modern methods to investigate item bias, specifically. In the psychometric literature, DIF analysis examines whether there is a systematic between-group difference in examinees’ probability to answer an item correctly while controlling for their ability [7]. While DIF is not equivalent to item bias, the former informs judgment on the latter when its presence is attributed to unintended content in the studied item [14, 15].
To examine systematic between-group differences in self-reported health survey data, we can apply existing DIF methods by generalizing the DIF concept with the following modifications:
@@
Some existing DIF methods examine both uniform and nonuniform DIF while many others do not. Uniform DIF occurs when respondents from one group consistently outperform respondents of the same latent trait from another group; nonuniform DIF exists when this performance difference is not consistent [16].
Most DIF methods currently used in health research, as well as in the broader DIF literature, are based on single-level modeling [e.g., 2]. They are not best suited for self-reported health survey data that has a nesting or hierarchical structure (e.g., individual responses are nested within larger units differentiated by population density, economy, and culture). These single-level DIF methods include the Mantel-Haenszel procedure [MH; 17], the logistic regression procedure [LR; 18], Lord’s Wald test [19], and the IRT likelihood-ratio test [IRT-LR; 20]. The first two do not require the IRT model and are labeled with “the observed-score approach,” whereas the latter two require the IRT model and are categorized as “the IRT approach” [21, 22]. An extensive review of these methods can be found in Chen [23].
Single-level modeling approaches have at least three problems when the response data has a hierarchical structure. First, they waste information and render inaccurate estimates because of the lack of consideration of variances between hierarchies of the data [24]. Second, they do not help model the effect of higher-level variables associated with DIF on the outcome variable at the lower level [25]. Third, their model estimates are limited to describing the sample and are not generalizable to describe the population [26]. Moreover, the single-level modeling for DIF analysis is limited to the analysis of two groups at a time, not suitable for DIF analyses involving more than two groups [23].
@@ 3 Solution: Multilevel modeling for DIF analysis
For the self-reported health survey data that is hierarchical in nature, multilevel modeling has been demonstrated to be a superior alternative to single-level modeling. It effectively addresses issues that arise in differential item functioning (DIF) methods built on single-level models. First, multilevel modeling is advantageous when fundamental assumptions in single-level modeling such as independence and homoscedasticity are violated, as is often the case with hierarchical data [24]. In this regard, multilevel modeling can account for correlated errors resulting from the dependency of individual observations within a nested structure, and it does not require the same sample size or variance for higher-level units due to its randomization approach [27].
@@
Second, multilevel modeling allows for the assessment of higher-level variables’ impact on DIF [25]. For instance, it can evaluate how cluster-level factors, like access to healthcare facilities, influence individual responses in health surveys, and identify whether these cluster characteristics contribute to DIF observed between population groups at the lower level of the model. Third, results from multilevel modeling are generalizable to the population because it incorporates random effects from higher-level clusters [26]. This ensures that conclusions about DIF in self-reported response data are applicable to a broader range of regions. Fourth, multilevel modeling accommodates multi-group analyses by using multiple dummy-coded variables to cover more than two groups [28]. For instance, it can include multiple age groups in health data, allowing researchers to analyze and compare health outcomes across various age categories while accounting for cluster-specific random effects.
@@
The first published work on DIF using multilevel modeling was by Swanson et al. [28], who adopted a multilevel logistic regression model. This approach significantly improved the accuracy of DIF estimates and pooled information across items to explain sources of DIF, when including random item effects. Subsequently, den Noortgate and de Boeck [29] proposed a series of multilevel models for DIF analysis, grounded in the Rasch model within item response theory (IRT). These models demonstrated the potential to integrate random effects associated with respondents, items, or other factors, along with item- or group-related covariates, to explain DIF. French and Finch [30] proposed a multilevel MH statistic that adjusts the traditional MH statistic using a ratio of random variances from the multilevel model from Swanson et al. [28]. Huang and Valdivia [31] introduced a multilevel Wald test for polytomous items based on Lord’s Wald $\chi^{2}$ test and a two-stage procedure about first identifying anchor items and then evaluating DIF items for non-anchor items.
In the following sections, we focus on the multilevel logistic regression procedure from Swanson et al. [28] because of its versatility. Compared to the models from den Noortgate and de Boeck [29] and from Huang and Valdivia [31], the multilevel logistic regression procedure does not require IRT or its stringent assumptions, such as local independence of item responses and monotonicity of the probability of the targeted response [32]. Meanwhile, to enhance the utility of the multilevel model in Swanson et al. [28], we have adapted it into a more general version that permits the inclusion of various types of random effects. For clarity, we also cover its single-level counterpart, the logistic regression procedure [18], to help readers understand how the multilevel logistic regression evolved from its origins in single-level modeling. Hereafter, we refer to the multilevel logistic regression procedure as “multilevel LR” and the logistic regression procedure as “traditional LR.” Also, we use the word “group” when referring to the group membership variable for group comparisons in DIF analysis, and the word “cluster” when referring to units of analyses across levels in the hierarchical data structure in multilevel modeling.
3.1 Multilevel LR
@@
The multilevel LR [28] was initially designed to include random effects associated with items, accounting for variations among items and offering more accurate DIF estimates compared to the traditional LR [18]. We have modified its notation to allow the incorporation of random effects from higher-level clusters in hierarchical data (e.g., access to healthcare, geographical locations, age groups, education levels). In this modified version, the level-1 submodel pertains to responses to a studied item from individual respondents (indexed by $p$) within a level-2 stratum (indexed by $i$), similar to the binary logistic regression model used in traditional LR but without an interaction term. This level-1 submodel is expressed as
$\mbox{logit}(\pi_{pi})=\beta_{0i}+\beta_{1i}\theta_{p}+\beta_{2i}g_{p},$ (1)
@@
$\beta_{2i}$ is the effect of respondents being in the reference group on the log-odds of the “probability targeted;”
@@
$g_{p}$ is a dummy variable representing the group membership of respondent $p$ ($g_{p}=0$ for the reference group, $g_{p}=1$ for each focal group).
The level-2 submodel in the multilevel LR accounts for random effects associated with higher-level clusters, written as
$\begin{array}[]{lll}\beta_{0i}&=&\delta_{00}+U_{0i},\\ \beta_{1i}&=&\delta_{10}+U_{1i},\\ \beta_{2i}&=&\delta_{20}+\delta_{21}v_{1}+\delta_{22}v_{2}+\cdots+\delta_{2V}v_{V}+U_{2i},\end{array}$
where
@@
$\delta_{20}$ is the average effect associated with being in the focal group;
$\delta_{21}$, $\delta_{22}$, $\cdots$, $\delta_{2V}$ denote the effect of covariates $v_{1}$, $v_{2}$, $\cdots$, $v_{V}$ (either continuous or dummy-coded) that are associated with level-2 clusters; they are useful to explain why DIF occurs, but can be omitted if no level-2 covariate is investigated here;
@@
$U_{1i}$ and $U_{2i}$ denote the random effect associated with level-2 clusters on the slopes $\beta_{1i}$ and $\beta_{2i}$, respectively;
@@
Using this multilevel LR, we can conclude that the studied item exhibits DIF if $\beta_{2i}$ is significantly different from 0.
3.2 Traditional LR
In the traditioinal LR proposed by Swaminathan and Rogers [18], the binary logistic regression model used for DIF analysis specifies the effect of respondents’ latent trait and group membership on the log-odds of the probability of the targeted response to a studied item. In Swaminathan and Rogers [18], the authors first rewrote the DIF model inherent in the widely used Mantel-Haenszel procedure using binary logistic regression:
$\mbox{logit}(\pi)=\beta_{0}+\sum_{k=1}^{K}\beta_{k}\theta_{k}+\tau g,$ (2)
@@ 3.2 Traditional LR
$\pi$ is the probability to provide the targeted response to the studied item;
$\beta_{0}$ denotes the average log-odds of the “probability targeted” in the reference group;
$\theta_{k}$ is the $k$th total score (in discrete units) in a fixed-item instrument for respondents;
$\tau$ is equivalent to the common odds ratio in the Mantel-Haenszel procedure;
$g$ is the group membership ($g$ = 0 for the reference group, $g$ = 1 for the focal group) of respondents.
Inspired by Mellenbergh [16], Swaminathan and Rogers [18] revised (2) by specifying separate terms for uniform and nonuniform DIF, written as
$\mbox{logit}(\pi)=\beta_{0}+\beta_{1}\theta+\beta_{2}g+\beta_{3}(\theta g),$ (3)
where
$\theta$ denotes respondents’ latent trait (proxied by the examinee’s total score);
$\beta_{1}$ is the effect of respondents’ latent trait on the log-odds of the “probability targeted” in the reference group;
$\beta_{2}$ is the effect of respondents being in the reference group on the log-odds of the “probability targeted;”
$\beta_{3}$ is the ability-moderated group difference in the effect on the log-odds of the “probability targeted;”
$g$ is respondents’ group membership ($g$ = 0 for the reference group, $g$ = 1 for the focal group).
Using the traditioinal LR specified in (3), we can conclude that the studied item exhibits nonuniform DIF if $\beta_{3}$ is significantly different from 0, and it exhibits uniform DIF if $\beta_{3}$ is not significantly different from 0 whereas $\beta_{2}$ is.
The traditional LR was established for analyzing two groups at a time. While a common practice in psychometrics is to perform pairwise comparisons using the traditional LR [e.g., 33], this approach can be computationally intensive and lead to increased Type-I error rates and reduced power [34, 35]. In contrast, the multilevel LR allows for multi-group analyses by employing multiple dummy variables to represent more than two groups [28], while controlling for cluster-specific random effects to ensure accuracy and generalizability.
@@ 4 Example: Analysis of population-density DIF in NSDUH data
To show the impact of multilevel modeling on DIF results, below we describe our DIF analyses when using both the multilevel LR and the traditional LR for comparison. The NSDUH data released in 2022 was utilized in these analyses. It provides comprehensive information on substance use and mental health trends among the U.S. population aged 12 and older. We are interested in examining DIF associated with population density because existing research has demonstrated that population density can affect the quality of life, health perceptions, and reported behaviors [36, 37, 38, 39]. In our multilevel analysis, we utilize a two-level model in which the level-2 clusters are defined by respondents’ education levels, as existing research has unveiled that an additional year of education could reduce the likelihood of reporting depression and anxiety [40].
Specifically, we want to examine whether DIF exists for respondents from high and low population density areas in the probability to answer “Yes” to a selected NSDUH item. This item asks respondents the following question: “Have you ever in your life had a period of time lasting several days or longer when most of the day you felt sad, empty or depressed?” The latent trait variable to be controlled for is a score assigned to respondents that reflects respondents’ psychological distress level over the past 30 days, based on their responses to another six questions. The reference group is the respondents from areas with low population density, and the focal group is the respondents from areas with high population density. The variable describing respondents’ education level is utilized for accounting for variations between the higher-level clusters. We acknowledge the potential for DIF in this latent trait variable originally derived from six survey items, but we do not examine DIF for this variable to focus on demonstrating the application of the multilevel LR rather than conducting a comprehensive DIF analysis across all measures.

### v2

# Examining Differential Item Functioning in Self-Reported Health Survey Data: Via Multilevel Modeling

@@ Abstract
Few health-related constructs or measures have received a critical evaluation in terms of measurement equivalence, such as self-reported health survey data. Differential item functioning (DIF) analysis is crucial for evaluating measurement equivalence in self-reported health surveys, which are often hierarchical in structure. Traditional single-level DIF methods in this case fall short, making multilevel models a better alternative. We highlight the benefits of multilevel modeling for DIF analysis, when applying a health survey data set to multilevel binary logistic regression (for analyzing binary response data) and multilevel multinominal logistic regression (for analyzing polytomous response data), and comparing them with their single-level counterparts. Our findings show that multilevel models fit better and explain more variance than single-level models. This article is expected to raise awareness of multilevel modeling and help healthcare researchers and practitioners understand the use of multilevel modeling for DIF analysis.
@@ 1 Introduction
To examine the disproportionate impact of health interventions or other determinants on health outcomes for certain populations, including disadvantaged minorities, it is crucial to establish measurement equivalence when evaluating outcomes across groups differing in gender, ethnicity, and other demographic factors. While validating psychometric instruments has been a routine practice in research, critical evaluations of measurement equivalence, particularly through differential item functioning (DIF) analysis, have historically been limited [1]. In recent years, however, there has been significant progress in this area, with a growing body of research employing DIF analysis to investigate measurement equivalence [e.g., 2, 3, 4, 5, 6, 7].
This growing research underscores the need for continued evaluation of health constructs and measures to ensure fairness and validity in assessing health outcomes for diverse populations. It is worth noting that some health studies did not clearly describe the method used for DIF analysis [e.g., 3], and conventional DIF methods based on single-level models were frequently used to analyze health survey data with a nesting structure [e.g., 2], for which recent DIF methods based on multilevel models are better alternatives in terms of accuracy of DIF estimation. By employing advanced methodologies for DIF, researchers can better evaluate health measures to ensure fairness and validity in assessing health outcomes for diverse populations.
@@ 2 Challenge: Examine DIF in self-reported health survey data
Existing research has shown that bias can manifest itself as construct bias, method bias, and item bias [e.g., 11, 12]. The framework of differential item functioning [DIF; 13] underpins many modern methods to investigate item bias, specifically. In the psychometric literature, DIF analysis examines whether there is a systematic between-group difference in examinees’ probability to answer an item correctly when controlling for their ability [14]. While DIF is not equivalent to item bias, the former can provide evidence for the latter when it arises from content or characteristics of the item that are unrelated to the intended construct being measured [15, 16]. Some existing DIF methods examine both uniform and nonuniform DIF while many others do not. Uniform DIF occurs when respondents from one group consistently outperform respondents of the same latent trait from another group; nonuniform DIF exists when this performance difference is not consistent [17].
To examine systematic between-group differences in self-reported health survey data, we can apply existing DIF methods by generalizing DIF-related concepts via the following modifications:
@@
The corresponding terminology may shift in the applications of DIF to address the unique contexts in health research.
Most DIF methods currently used in health research, as well as in the broader DIF literature, are based on single-level modeling [e.g., 2]. They are not best suited for self-reported health survey data that has a nesting or hierarchical structure (e.g., individual responses are nested within larger units differentiated by population density, economy, and culture). These single-level DIF methods include the logistic regression procedure [LR; 18], the Mantel-Haenszel procedure [MH; 19], Lord’s Wald test [20], and the IRT likelihood-ratio test [IRT-LR; 21]. The first two do not require the IRT model and are labeled with “the observed-score approach,” whereas the latter two require the IRT model and are categorized as “the IRT approach” [22, 23]. An extensive review of these methods can be found in Chen [24].
Single-level modeling approaches have at least three problems when the response data has a hierarchical structure. First, they waste information and render inaccurate estimates because of the lack of consideration of variances between hierarchies of the data [25]. Second, they do not help model the effect of higher-level variables associated with DIF on the outcome variable at the lower level [26]. Third, their model estimates are limited to describing the sample and are not generalizable to describe the population [27]. Moreover, most of the single-level DIF methods are limited to the analysis of two groups at a time, showing inflated error rates for DIF analyses involving more than two groups [24, 28, 29].
@@ 3 Solution: Multilevel modeling for DIF analysis
For the self-reported health survey data that is hierarchical in nature, multilevel modeling has been demonstrated to be a superior alternative to single-level modeling. It effectively addresses issues that arise in DIF methods built on single-level models. First, multilevel modeling is advantageous when fundamental assumptions in single-level modeling such as independence and homoscedasticity are violated, as is often the case with hierarchical data [25]. In this regard, multilevel modeling can account for correlated errors resulting from the dependency of individual observations within a nested structure, and it does not require the same sample size or variance for higher-level units due to its randomization approach [30].
@@
Second, multilevel modeling allows for the assessment of higher-level variables’ impact on DIF [26]. For instance, it can evaluate how cluster-level factors, like access to healthcare facilities, influence individual responses in health surveys, and identify whether these cluster characteristics contribute to DIF observed between population groups (e.g., gender, ethnicity, location) at the lower level of the model. Third, results from multilevel modeling are generalizable to the population because it incorporates random effects from higher-level clusters [27]. This ensures that conclusions about DIF in self-reported response data are applicable to a broader range of demographic groups. Fourth, multilevel modeling accommodates multi-group analyses by using multiple dummy-coded variables to cover more than two groups [31]. For instance, it can include multiple age groups in health data, allowing researchers to analyze and compare health outcomes across various age categories while accounting for cluster-specific random effects.
@@
The first published work on DIF using multilevel modeling was by Swanson et al. [31], who adopted a multilevel binary LR model. This approach significantly improved the accuracy of DIF estimates and pooled information across items to explain sources of DIF, when including random item effects. Subsequently, den Noortgate and de Boeck [32] proposed a series of multilevel models for DIF analysis, grounded in the Rasch model within item response theory (IRT). These models demonstrated the potential to integrate random effects associated with respondents, items, or other factors, along with item- or group-related covariates, to explain DIF. French and Finch [33] proposed a multilevel MH statistic that adjusts the traditional MH statistic using a ratio of random variances from the multilevel model from Swanson et al. [31]. Huang and Valdivia [34] introduced a multilevel Wald test for polytomous items based on Lord’s Wald $\chi^{2}$ test.
Each of these multilevel approaches offers unique benefits, but they also come with specific limitations. For instance, the models from den Noortgate and de Boeck [32] leverage the rigor of IRT, allowing for a deeper integration of item- and respondent-level characteristics and providing insights into hierarchical data structures. However, these models rely on assumptions such as local independence of item responses and monotonicity of response probabilities, which can be restrictive or challenging to meet in certain applications [35]. In contrast, the multilevel LR model adapted from Swanson et al. [31] offers greater versatility by not requiring such stringent assumptions, making it suitable for contexts where IRT assumptions might not hold.
The choice of method depends on the research objectives and the nature of the data. For example, the IRT-based multilevel models proposed by den Noortgate and de Boeck [32] are particularly useful when the goal is to investigate complex random effects structures or when there is a need for a model grounded in the theoretical framework of IRT. However, these models may be computationally intensive and less intuitive for practitioners unfamiliar with IRT. Conversely, the multilevel LR model provides a balance of flexibility and interpretability, making it a practical choice for broader applications.
In the following sections, we demonstrate the use of two types of multilevel models corresponding to binary and polytomous response data, based on Swanson et al. [31] because of the versatility. First, for analyzing DIF in binary response data from a survey item, we utilize the multilevel binary LR adapted from Swanson et al. [31]. Its single-level counterpart, the LR procedure [18], is reviewed in Appendix A, hereafter referred to as “traditional LR”. Second, for analyzing DIF in polytomous response data, we propose the multilevel multinominal LR. Its single-level counterpart is the multinominal LR, a type of multicategory logit models for unordered categories and is seen as a generalization of the binary LR [36]. Throughout this writing, we use the word “group” when referring to the group membership variable for group comparisons in DIF analysis, and the word “cluster” when referring to units of analyses across levels in the hierarchical data structure in multilevel modeling.
3.1 Multilevel Binary LR for Binary Responses
@@
The multilevel binary LR [31] was initially designed to include random effects associated with items, accounting for variations among items and offering more accurate DIF estimates compared to the traditional LR [18]. We have modified its notation to allow the incorporation of more than one focal group and of random effects from higher-level clusters in hierarchical data (e.g., access to healthcare, geographical locations, age groups, education levels).
In this modified version, the level-1 submodel pertains to responses to a studied item from individual respondents, indexed by $p$, within a level-2 stratum, indexed by $i$. This two-level binary LR model is specified as follows:
Level $\begin{array}[]{lrll}\textbf{Level 1:}&&&\\ &\mbox{logit}(\pi_{pi})&=&\mbox{log}(\dfrac{\pi_{pi}}{1-\pi_{pi}})=\beta_{0i}+\beta_{1i}\theta_{p}+\boldsymbol{\beta_{2i}g_{p}},\\ \textbf{Level 2:}&&&\\ &\beta_{0i}&=&\delta_{00}+U_{0i},\\ &\beta_{1i}&=&\delta_{10}+U_{1i},\\ &\boldsymbol{\beta_{2i}}&=&\boldsymbol{\delta_{20}}+U_{2i}.\end{array}$ (1)
@@
$\boldsymbol{\beta_{2i}}$ is a vector for the deviation effect in a focal group from the reference group;
@@
$\boldsymbol{g_{p}}$ is a vector containing dummy-coded variables representing the group membership of respondent $p$, specified as $\begin{bmatrix}g_{p1}&g_{p2}&\cdots&g_{pJ}\end{bmatrix}$ ($g_{pj}=1$ for being in the target group, indexed by $j=1,2,\cdots,J$);
@@
$\boldsymbol{\delta_{20}}$ is a vector for the average effect associated with being in each focal group;
@@
$U_{1i}$ and $U_{2i}$ denote the random effect associated with level-2 clusters on the slopes $\beta_{1i}$ and $\boldsymbol{\beta_{2i}}$, respectively;
@@
The level-1 submodel of the multilevel binary LR is a binary LR model, similar to the one in traditional LR but without an ability-group interaction term. It is limited to binary response data analysis. The level-2 submodel in the multilevel binary LR accounts for random effects associated with higher-level clusters. When using this multilevel binary LR, the studied item exhibits DIF if any coefficient value in the vector $\boldsymbol{\beta_{2i}}$ is significantly different from 0.
3.2 Multilevel Multinominal LR for Polytomous Responses
We propose the multilevel multinominal LR for analyzing DIF in polytomous response data. Its level-1 submodel is the multinominal extension of the traditional LR but without the ability-group interation term. Its level-2 submodel incorporates random effects tied to higher-level clusters. This two-level multilevel multinominal LR is written as
Level $\begin{array}[]{lrll}\textbf{Level 1:}&&&\\ &\mbox{log}(\dfrac{\pi_{c.pi}}{\pi_{B.pi}})&=&\beta_{0i}+\beta_{1i}\theta_{p}+\boldsymbol{\beta_{2i}g_{p}},\\ \textbf{Level 2:}&&&\\ &\beta_{0i}&=&\delta_{00}+U_{0i},\\ &\beta_{1i}&=&\delta_{10}+U_{1i},\\ &\boldsymbol{\beta_{2i}}&=&\boldsymbol{\delta_{20}}+U_{2i}.\end{array}$ (2)
@@ 3.2 Multilevel Multinominal LR for Polytomous Responses
$\pi_{c.pi}$ is the probability for respondent $p$ to provide a non-baseline response, indexed by $c$, to the studied item ($c=1,2,\cdots,C$);
$\pi_{B.pi}$ is the probability for respondent $p$ to provide the baseline response, denoted by $B$, to the studied item.
The level-1 submodel pairs a selected baseline response with each of the other responses to the survey item, and it has $C$ equations, corresponding to the number of non-baseline responses. Parameters for these equations are estimated separately. For example, when the response data has three unique responses (e.g., 0, 1, and 2) and we choose one as the baseline (e.g., 0), $C=2$ and the level-1 submodel is written into
Equation $\begin{array}[]{lrll}\text{Equation 1:}&\mbox{log}(\dfrac{\pi_{1.pi}}{\pi_{B.pi}})&=&\beta_{1.0i}+\beta_{1.1i}\theta_{p}+\boldsymbol{\beta_{1.2i}g_{p}},\\ \text{Equation 2:}&\mbox{log}(\dfrac{\pi_{2.pi}}{\pi_{B.pi}})&=&\beta_{2.0i}+\beta_{2.1i}\theta_{p}+\boldsymbol{\beta_{2.2i}g_{p}}.\\ \end{array}$
@@ 4 Example: Analysis of population-density DIF in NSDUH data
Existing research has demonstrated that population density can affect the quality of life and health perceptions [37, 38, 39, 40]. We utilize the NSDUH data released in 2022 to examine population-density-related DIF in the self-reported health survey questions evaluating depression. This data provides comprehensive information on substance use and mental health trends among the U.S. population aged 12 and older. To show the impact of multilevel modeling on DIF results, below we compare the results from the multilevel binary LR and the multilevel multinominal LR with their single-level counterparts. We use respondents’ education levels as the level-2 clusters in multilevel modeling, for existing research has unveiled that an additional year of education could reduce the likelihood of reporting depression and anxiety [41].
Specifically, we want to examine whether DIF exists for respondents from high and low population density areas in the probability to answer “Yes” to a selected NSDUH survey item. Hence, the targeted response in this DIF analysis is “Yes”. This item asks respondents the following question: “Have you ever in your life had a period of time lasting several days or longer when most of the day you felt sad, empty or depressed?” The latent trait variable to be controlled for is respondents’ psychological distress level over the past 30 days, proxied by a score based on their responses to another six survey questions. We acknowledge the potential for DIF in this derived score; however, we do not examine DIF in this instance to maintain our focus on demonstrating the application of multilevel modeling. The reference group is the respondents from areas with low population density, and there are two focal groups: one is the respondents from areas with high population density, and the other is from non-classified areas.